OpenStudy (anonymous):

proportions question: 7/21=1/n

OpenStudy (anonymous):

 Multiply diagonally then divide horizontally to get your answer

OpenStudy (anonymous):

1/3?

OpenStudy (blurbendy):

7 / 21 = 1 / n 21 = 7n n = 3

OpenStudy (anonymous):

 cross multiply then you will get what @blurbendy got :) (sorry for bad drawing xD)